# Schedule On the Fly: Diffusion Time Prediction for Faster and Better Image Generation

### v1

@@ 2.2 Noise Schedule
There exist some methods fine-tuning the scheduler to speed up sampling or improve image quality. Xia et al. [52] predict a new denoising time for each step to find a more accurate integration direction, Sabour et al. [40] uses stochastic calculus to find the optimal sampling plan for different schedulers and different models. Wang et al. [49] leverages reinforcement learning to automatically search for an optimal sampling scheduler. Some one-step generators[29, 30, 13] with diffusion distillation [26] also achieve impressive performance.
@@ 2.3 Reinforcement Learning and Learning from Human Feedback
Reinforcement Learning from Human Feedback (RLHF) has recently gained significant attention in the field of large language models (LLMs) [42, 1, 47] and is gradually expanding into other domains. Advances in diffusion models have increasingly incorporated reward models, trained on human preferences, to enhance alignment with human values [21, 54, 51]. Some recent works have also studied the RLHF for few-step generative models [27, 32]. However, most approaches only employ reward models as substitutes for diffusion loss or in conjunction with diffusion processes to align with human preferences, which significantly differs from the Proximal Policy Optimization (PPO) paradigm [42] commonly used in RLHF for LLMs. This divergence arises primarily because the multi-step denoising process in DPMs does not naturally yield likelihoods in the same manner as language models, posing challenges for the direct application of PPO. In this paper, however, we introduce an approach that, to the best of our knowledge, is the first to treat each denoising step as a distribution analogous to an action in reinforcement learning. This enables the use of well-established optimization techniques from LLM-based RLHF within the context of diffusion models.
@@ 4.1 Implementation details
We use the AdamW optimizer with $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.99, a constant learning rate of 1e-5, and a maximum gradient norm of 1.0. Our TPM module typically only requires around 200 training steps with a batch size of 256.
@@ 4.3 Quantitative Results
Moreover, the metrics representing human preference improve the most. For example, by generating with just 15.28 steps on average, TPDM-SD3-Medium gets a 29.59 HPS score, +1.07 higher than Stable Diffusion 3 with similar steps, also +0.47 higher than the original 28-step results. This may attributes to the reward model we utilized to calculate advantage in optimization, leading to aesthetically pleasing images aligned with human preferences.

### v2

@@ 2.2 Noise Schedule
There exist some methods fine-tuning the scheduler to speed up sampling or improve image quality. Xia et al. [60] predict a new denoising time for each step to find a more accurate integration direction, Sabour et al. [46] uses stochastic calculus to find the optimal sampling plan for different schedulers and different models. Wang et al. [56] leverages reinforcement learning to automatically search for an optimal sampling scheduler. Also, adjust the schedule is also effective is also effective in other areas such as molecular generation [17]. Some one-step generators[34, 35, 16] with diffusion distillation [31] also achieve impressive performance.
@@ 2.3 Reinforcement Learning and Learning from Human Feedback
Reinforcement Learning from Human Feedback (RLHF) has recently gained significant attention in the field of large language models (LLMs) [48, 1, 53] and is gradually expanding into other domains. Advances in diffusion models have increasingly incorporated reward models to enhance alignment with human preferences [25, 62, 59]. By treating latents as actions, diffusion models can be optimized with policy gradient (DDPO [3], DPOK[12]), actor-critic framework (DACER [58]), direct preference optimization (Diffusion-DPO [55]) and other reinforcement algorithms. Some recent works have also studied the RLHF for few-step generative models [32, 37]. Since these approaches rely on the Gaussian reverse processes of the diffusion SDE sampler, however, they are challenging to apply in flow models. On the contrary, in this paper we treat the diffusion time as the action instead of latents. Our goal is to adjust the schedule to achieve better quality with fewer steps, offering a more general and flexible solution.
@@ 4.1 Implementation details
We use the AdamW optimizer with $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.99, a constant learning rate of 1e-5, and a maximum gradient norm of 1.0. Our TPM module typically requires only 200 sampled trajectories, 4 PPO epochs, and a batch size of 256 to effectively update its parameters.
@@ 4.3 Quantitative Results
Moreover, the metrics representing human preference improve the most. For example, by generating with just 15.28 steps on average, TPDM-SD3-Medium gets a 29.59 HPS score, +1.07 higher than Stable Diffusion 3 with similar steps, also +0.47 higher than the original 28-step results. This may attributes to the reward model we utilized to calculate advantage in optimization, leading to aesthetically pleasing images aligned with human preferences. Furthermore, since other scheduling-tuning methods only conduct experiments on classical diffusion models such as Stable Diffusion 1.5 [45], we conduct an apple-to-apple comparison with these algorithms as well, which is demonstrated in Appendix Table 4.
@@ B.1 Figure 1
Method Inference Steps FID CLIP-T Aesthetic v2 Pick Score HPSv2.1
DPM-Solver++ [30] 10 21.08 0.314 5.219 21.30 24.00
DPM-Solver++ & AYS [46] 10 18.81 0.313 5.156 21.31 24.40
DPM-Solver++ & GITS [7] 10 18.32 0.313 5.108 21.19 23.72
DPM-Solver++ & TPM 9.890 18.31 0.313 5.118 21.32 24.29
Table 4: Experiments on Stable Diffusion v1.5 [45]. 9.890 is the average step TPM used.
Appendix C Applicaion on Denoising Diffusion Probabilistic Models (DDPMs)
Prior to the widespread use of flow models, diffusion models dominates image generation, such as DDPM [13] and LDM [45]. They demonstrate several difference from flow models during inference. First, they directly predict the noise ($\epsilon$) added to the data at each step, rather than predicting the velocity vector field. Second, both the diffusion time $t$ and the noise level $\sigma_{t}$ are typically discrete, unlike continuous time in flow models. Moreover, the noise level $\sigma_{t}$ does not exhibit a simple linear relationship with the diffusion time $t$.
In this section, we integrate TPM into Stable Diffusion 1.5 [45], demonstrating that TPM also works in DDPMs. When determining the next diffusion time, TPM predicts the Beta distribution as in flow models, and then quantizes it to the closest $t$ in all $T$ discrete steps, thereby obtaining the corresponding noise level $\sigma_{t}$.
We can generally express their relationship as follows:
$\beta_{t}=\left(\sqrt{\beta_{start}}+\frac{t}{T-1}\left(\sqrt{\beta_{end}}-\sqrt{\beta_{start}}\right)\right)^{2},$ (10)
$\alpha_{t}=1-\beta_{t},$ (11)
$\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s},$ (12)
$\sigma_{t}=\sqrt{\frac{1-\bar{\alpha}_{t}}{\bar{\alpha}_{t}}},$ (13)
where $\beta_{t}$ is the variance of the noise added at diffusion time $t$. $\alpha_{t}$ is the proportion of the original signal remaining. $\sigma_{t}$ is the total noise level, and $T$ is the total number of diffusion steps while training, typically set to 1000. $\beta_{start}$ and $\beta_{end}$ are both hyper-parameter, controlling the initial and final noise variances.
We compare the performance of TPM with other scheduled fine-tuning methods applied to diffusion models. The evaluation metrics are calculated using 5,000 prompts from the COCO 2017 validation set, adhering to the evaluation protocol outlined in Table 1. Notably, our findings reveal that TPM can be effectively combined with higher-order solvers, such as DPM-Solver++ [30], resulting in a significant improvement of -2.79 in FID, surpassing the performance of other methods. It is important to note that GITS [7] requires access to the statistics of the evaluation set. While this approach is advantageous for large-scale image generation, it may offer limited benefits when generating a single image. Despite that GITS makes use of additional information, TPM consistently outperforms GITS across all evaluation metrics.